OpenStudy (anonymous):

Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form.

OpenStudy (nubeer):

m = 2 x1 = -3 , y = 4 now use this equation y-y1 = m(x-x1) plug in the given data u will get the answer

OpenStudy (anonymous):

use y-y1=m(x-x1) if m=slope and (x1,y1) are the point given :)